# Moser’s twist theorem revisited

### v2

@@ 5. Verification of Criterion 3 for all $\kappa\geq\kappa_{0}$
    (40) $\tilde{\mathbf{K}}_{r}^{1}(\kappa,r)\leq C_{1}q_{r}^{-\varepsilon/3}$
@@
We prove Lemma 5.1 by induction. At first, we check the initial step for $\kappa=\kappa_{0}$. Let $v\in(\kappa_{0})_{II}$. Let us recall $s(v)=\frac{r(v)}{\Theta(v)}$. Assume $v$ connects $(x_{i},y_{i})$ and $(x_{j},y_{j})$. From Definition 3.3 and Lemma 2.2, there exists a constant $C_{3}>0$ such that
$|r(v)|=|y_{j}-y_{i}|\leq C_{3}q_{n_{0}}^{-2-\frac{\varepsilon}{2}}.$
By Lemma 3.6, we have $\Theta(v)\sim q_{n_{0}}^{-1}$. It follows from (16) that for $(v_{1},v_{2})\in(\kappa_{0},r)_{II}$ and $n_{0}$ large enough
(45) $\nabla^{1}(v_{1},v_{2})=\frac{s(v_{2})-s(v_{1})}{\Theta(v_{1},v_{2})}\leq\frac{s(v_{2})-s(v_{1})}{\Theta(v_{1})}\leq C_{3}q_{n_{0}}^{-\frac{\varepsilon}{2}}\leq\frac{\varepsilon}{1920A_{\alpha}}.$
@@
$\lambda(v_{1},\ldots,v_{4})\sim\Theta(v_{1},\ldots,v_{4}).$
@@
$\nabla^{2}(v_{1},\ldots,v_{4})=\frac{\Theta(v_{1},\ldots,v_{4})}{\lambda(v_{1},\ldots,v_{4})}(\nabla^{1}(v_{3},v_{4})-\nabla^{1}(v_{1},v_{2}))\lesssim q_{n_{0}}^{-\frac{\varepsilon}{2}}.$
Note that $\tilde{N}(\kappa)\leq n_{\kappa}$ for each $\kappa\in\mathbb{N}$. It follows from (24) that $\nabla^{1}(\kappa_{0},\tilde{N}(\kappa_{0}))\leq\nabla^{1}(\kappa_{0},n_{0})$. Hence, (43) and (44) hold for $\kappa=\kappa_{0}$.
@@ 5.1. Recursive property
Let $G(v):=\frac{\Theta(v)}{\lambda(v)}$. Given $\kappa\in\mathbb{N}$, if we assume
@@
$|G(v_{2})-G(v_{1})|\leq C_{4}q_{\tilde{N}}^{-\frac{\varepsilon}{3}}.$
@@
Proof   By the arithmetics of $\alpha$, one can find an integer $M>0$ independent of $\kappa$ such that
@@
$|\Theta(v_{2})-\Theta(v_{1})|\leq|\Theta(F^{Mq_{\tilde{N}}}(v_{1})-\Theta(v_{1})|.$
@@
$\mathbb{K}_{\tilde{N}}^{0}(\kappa)_{II}\leq C_{1}q_{\tilde{N}}^{-\frac{\varepsilon}{3}}.$
@@
$\exp\left(-C_{1}q_{\tilde{N}}^{-\frac{\varepsilon}{3}}\right)\leq\frac{\Theta(F^{q_{\tilde{N}}}(v))}{\Theta(v)}\leq\exp\left(C_{1}q_{\tilde{N}}^{-\frac{\varepsilon}{3}}\right).$
@@
$\exp\left(-MC_{1}q_{\tilde{N}}^{-\frac{\varepsilon}{3}}\right)\leq\frac{\Theta(F^{Mq_{\tilde{N}}}(v_{1})}{\Theta(v_{1})}\leq\exp\left(MC_{1}q_{\tilde{N}}^{-\frac{\varepsilon}{3}}\right).$
@@
$\frac{|\Theta(F^{Mq_{\tilde{N}}}(v_{1})-\Theta(v_{1})|}{\Theta(v_{1})}\leq\max\left\{\exp\left(MC_{1}q_{\tilde{N}}^{-\frac{\varepsilon}{3}}\right)-1,1-\exp\left(-MC_{1}q_{\tilde{N}}^{-\frac{\varepsilon}{3}}\right)\right\},$
@@
$|\Theta(F^{Mq_{\tilde{N}}}(v_{1})-\Theta(v_{1})|\leq 2MC_{1}q_{\tilde{N}}^{-\frac{\varepsilon}{3}}\Theta(v_{1}).$
@@
$|G(v_{2})-G(v_{1})|\leq\frac{|\Theta(F^{Mq_{\tilde{N}}}(v_{1})-\Theta(v_{1})|}{\lambda(v_{1})}\leq 2MC_{1}q_{\tilde{N}}^{-\frac{\varepsilon}{3}}\frac{\Theta(v_{1})}{\lambda(v_{1})}$
[…]
Let $\beta=\frac{\varepsilon}{3}$. Based on [11, Corollary 3.1], we have an improved estimate
(46) $\nabla^{2}(\kappa,\tilde{N},\tilde{N})\leq C_{0}q_{\tilde{N}}^{-\eta-\beta},$
where $\eta:=2^{-10}\beta$. By using Lemma 5.3 and (46), we obtain the following recursive property.
@@
$|G(v_{2})-G(v_{1})|\leq C_{4}q_{\tilde{N}}^{-\frac{\varepsilon}{3}}.$
@@
Proof   Let $v\in(\kappa+1)_{II}$ and $v_{1}$ be a contiguous vector in $(\kappa)_{II}$, satisfying $\lambda(v)\sim\frac{1}{2}\lambda(v_{1})$ as $\kappa\to+\infty$. It follows from the estimate $|G(v_{2})-G(v_{1})|\leq C_{4}q_{\tilde{N}}^{-\varepsilon/3}$ that $\Theta(v)\sim\frac{1}{2}\Theta(v_{1})$.
Let $v_{2}$ be a $\kappa$-vector such that $(v_{1},v_{2})\in(\kappa,\tilde{N}(\kappa))_{II}$, which consequently implies $(v+v_{1},v_{2})\in(\kappa,\tilde{N}(\kappa))_{II}$. From the convex combination
$s(v+v_{1})=\frac{\Theta(v)}{\Theta(v)+\Theta(v_{1})}s(v)+\frac{\Theta(v_{1})}{\Theta(v)+\Theta(v_{1})}s(v_{1}),$
we derive the decomposition of $\nabla^{1}(v+v_{1},v_{2})$:
$\nabla^{1}(v+v_{1},v_{2})$ $=\frac{s(v_{2})-bs(v)-b_{1}s(v_{1})}{\Theta(v+v_{1},v_{2})}$
$=\frac{\Theta(v,v_{2})}{\Theta(v+v_{1},v_{2})}b\nabla^{1}(v,v_{2})+\frac{\Theta(v_{1},v_{2})}{\Theta(v+v_{1},v_{2})}b_{1}\nabla^{1}(v_{1},v_{2})$
(50) $=c\nabla^{1}(v,v_{2})+c_{1}\nabla^{1}(v_{1},v_{2}),$
@@
$b=\frac{\Theta(v)}{\Theta(v)+\Theta(v_{1})},\quad b_{1}=\frac{\Theta(v_{1})}{\Theta(v)+\Theta(v_{1})},\quad c=\frac{\Theta(v,v_{2})}{\Theta(v+v_{1},v_{2})},\quad c_{1}=\frac{\Theta(v_{1},v_{2})}{\Theta(v+v_{1},v_{2})}.$
@@
This establishes that, up to errors bounded by $\nabla^{2}(\kappa,\tilde{N}(\kappa),\tilde{N}(\kappa))$, the $\nabla^{1}$ estimates (and similarly for $\nabla^{2}$) remain valid when replacing one vector in the set $(\kappa)_{II}$ with a vector in the set $(\kappa+1)_{II}$. By iterating this argument, we can successively replace all vectors in the set $(\kappa)_{II}$ with vectors in the set $(\kappa+1)_{II}$, thereby proving claims (47) and (48).
For the claim (49), consider $(v_{1},v_{2})\in(\kappa+1,\tilde{N}(\kappa+1))_{II}$ and let $v_{3}\in(\kappa+1)_{II}$ satisfy $\Theta(v_{1},v_{3})\sim q_{\tilde{N}(\kappa)}^{-1}$. Then $\nabla^{1}(v_{1},v_{2})$ admits the decomposition:
@@
$d_{1}=-\frac{\Theta(v_{1},v_{3})}{\Theta(v_{1},v_{2})},\quad d_{2}=\frac{\Theta(v_{2},v_{3})}{\Theta(v_{1},v_{2})},$
@@
$\nabla^{2}(\kappa,\tilde{N},\tilde{N})\leq C_{0}q_{\tilde{N}}^{-\eta-\beta}.$
@@
$\tilde{N}_{0}:=\tilde{N}(\kappa_{0}),\quad\varepsilon_{1}:=\frac{\varepsilon}{960A_{\alpha}},$
$C_{5}:=100(A_{\alpha}+1)^{3}C_{0},\quad C_{6}:=\frac{C_{5}\log_{2}(2A_{\alpha})}{1-\sqrt{2}^{-(\beta+\eta)}}.$
@@
For each $m>0$, if we assume that for each $\kappa\in[\kappa_{0},\kappa_{0}+m-1]$,
@@
Proof   By Lemma 5.3 and Lemma 5.4,
@@
$\leq\frac{\varepsilon_{1}}{2}+C_{5}q_{\tilde{N}_{0}}^{-\eta-\beta}.$
@@
$\begin{split}\nabla^{1}(\kappa_{0}+m,\tilde{N}(\kappa_{0}+m))\leq\frac{\varepsilon_{1}}{2}+C_{6}\sqrt{2}^{-(\beta+\eta)\kappa_{0}}.\end{split}$
@@
(52) $\kappa_{0}>\frac{2\log_{2}\frac{2C_{6}}{\varepsilon_{1}}}{\beta+\eta},$

### v3

@@ 5. Verification of Criterion 3 for all $\kappa\geq\kappa_{0}$
    (56) $\widetilde{\mathbf{K}}_{r}^{1}(\kappa,r)\leq C_{1}q_{r}^{-\varepsilon/3}$
@@
We prove Lemma 5.1 by induction. At first, we check the initial step for $\kappa=\kappa_{0}$. Denote $n_{0}:=n_{k_{0}}$ for simplicity and let $v\in(\kappa_{0})_{II}$. Let us recall
$\tilde{N}(k_{0})=n_{k_{0}}-1=n_{0}-1,\quad s(v)=\frac{r(v)}{\Theta_{1}(v)}.$
Assume $v$ connects $(x_{i},y_{i})$ and $(x_{j},y_{j})$. From Definition 3.4 and Lemma 2.2, there exists a constant $C_{3}>0$ such that
(61) $|r(v)|=|y_{j}-y_{i}|\leq C_{3}q_{n_{0}}^{-2-\varepsilon}.$
By Lemma 3.7, we have $\Theta_{1}(v)\sim q_{n_{0}}^{-1}$. It follows from (21) that for $(v_{1},v_{2})\in(\kappa_{0},r)_{II}$ and $n_{0}$ large enough
(62) $\nabla^{1}(v_{1},v_{2})=\frac{s(v_{2})-s(v_{1})}{\Theta_{2}(v_{1},v_{2})}\leq\frac{s(v_{2})-s(v_{1})}{\Theta_{1}(v_{1})}\leq C_{3}q_{n_{0}}^{-\varepsilon}\leq\frac{\varepsilon}{1920A_{\alpha}}.$
@@
$\lambda_{4}(v_{1},\ldots,v_{4})\sim\Theta_{4}(v_{1},\ldots,v_{4}).$
@@
$\nabla^{2}(v_{1},\ldots,v_{4})=\frac{\Theta_{4}(v_{1},\ldots,v_{4})}{\lambda_{4}(v_{1},\ldots,v_{4})}(\nabla^{1}(v_{3},v_{4})-\nabla^{1}(v_{1},v_{2}))\lesssim q_{n_{0}}^{-\varepsilon}\leq C_{0}q_{\tilde{N}(k_{0})}^{-\frac{\varepsilon}{3}}.$
Note that $\tilde{N}(\kappa)=n_{\kappa}-2\gamma_{0}\leq n_{\kappa}$ for each $\kappa\in\mathbb{N}$. It follows from Lemma 4.4 that
$\nabla^{1}(\kappa_{0},\tilde{N}(\kappa_{0}))\leq\nabla^{1}(\kappa_{0},n_{0}).$
Hence, (59) and (60) hold for $\kappa=\kappa_{0}$.
@@ 5.1. Recursive property
Let $\Phi(v):=\frac{\Theta_{1}(v)}{\lambda_{1}(v)}$. Given $\kappa\in\mathbb{N}$, if we assume
@@
$|\Phi(v_{2})-\Phi(v_{1})|\leq C_{4}q_{\tilde{N}}^{-\frac{\varepsilon}{3}}.$
@@
Proof.
By the arithmetics of $\alpha$, one can find an integer $M>0$ independent of $\kappa$ such that
@@
$|\Theta_{1}(v_{2})-\Theta_{1}(v_{1})|\leq|\Theta_{1}(F^{Mq_{\tilde{N}}}(v_{1})-\Theta_{1}(v_{1})|.$
@@
$\mathbb{K}_{\tilde{N}}^{0}(\kappa)\leq C_{1}q_{\tilde{N}}^{-\frac{\varepsilon}{3}}.$
@@
$\exp\left(-C_{1}q_{\tilde{N}}^{-\frac{\varepsilon}{3}}\right)\leq\frac{\Theta_{1}(F^{q_{\tilde{N}}}(v))}{\Theta_{1}(v)}\leq\exp\left(C_{1}q_{\tilde{N}}^{-\frac{\varepsilon}{3}}\right).$
@@
$\exp\left(-MC_{1}q_{\tilde{N}}^{-\frac{\varepsilon}{3}}\right)\leq\frac{\Theta_{1}(F^{Mq_{\tilde{N}}}(v_{1})}{\Theta_{1}(v_{1})}\leq\exp\left(MC_{1}q_{\tilde{N}}^{-\frac{\varepsilon}{3}}\right).$
@@
$\frac{|\Theta_{1}(F^{Mq_{\tilde{N}}}(v_{1})-\Theta_{1}(v_{1})|}{\Theta_{1}(v_{1})}\leq\max\left\{\exp\left(MC_{1}q_{\tilde{N}}^{-\frac{\varepsilon}{3}}\right)-1,1-\exp\left(-MC_{1}q_{\tilde{N}}^{-\frac{\varepsilon}{3}}\right)\right\},$
@@
$|\Theta_{1}(F^{Mq_{\tilde{N}}}(v_{1})-\Theta_{1}(v_{1})|\leq 2MC_{1}q_{\tilde{N}}^{-\frac{\varepsilon}{3}}\Theta_{1}(v_{1}).$
@@
$|\Phi(v_{2})-\Phi(v_{1})|\leq\frac{|\Theta_{1}(F^{Mq_{\tilde{N}}}(v_{1})-\Theta_{1}(v_{1})|}{\lambda(v_{1})}\leq 2MC_{1}q_{\tilde{N}}^{-\frac{\varepsilon}{3}}\frac{\Theta_{1}(v_{1})}{\lambda(v_{1})}$
By (58), we know that $\frac{\Theta_{1}(v_{1})}{\lambda(v_{1})}\leq C_{2}$. This completes the proof. ∎
Based on [11, Corollary 3.1], we have an improved estimate
(63) $\nabla^{2}(\kappa,\tilde{N},\tilde{N})\leq C_{0}q_{\tilde{N}}^{-\eta-\frac{\varepsilon}{3}},$
where $\eta:=\frac{1}{2^{10}}\frac{\varepsilon}{3}$. By using Lemma 5.3 and (63), we obtain the following recursive property.
@@
$|\Phi(v_{2})-\Phi(v_{1})|\leq C_{4}q_{\tilde{N}}^{-\frac{\varepsilon}{3}}.$
@@
Proof.
Let $v\in(\kappa+1)_{II}$ and $v_{1}$ be a contiguous chord in $(\kappa)_{II}$, satisfying $\lambda_{1}(v)\sim\frac{1}{2}\lambda(v_{1})$ as $\kappa\to+\infty$. It follows from the estimate $|\Phi(v_{2})-\Phi(v_{1})|\leq C_{4}q_{\tilde{N}}^{-\varepsilon/3}$ that $\Theta_{1}(v)\sim\frac{1}{2}\Theta_{1}(v_{1})$.
There exists a $\kappa$-chord $v_{2}$ such that $(v_{1},v_{2})\in(\kappa,\tilde{N}(\kappa))_{II}$, which consequently implies $(v+v_{1},v_{2})\in(\kappa,\tilde{N}(\kappa))_{II}$. Similar to the proof of Lemma 4.4, we have
(67) $\nabla^{1}(v+v_{1},v_{2})=c\nabla^{1}(v,v_{2})+c_{1}\nabla^{1}(v_{1},v_{2}),$
@@
$b=\frac{\Theta_{1}(v)}{\Theta_{1}(v)+\Theta_{1}(v_{1})},\quad b_{1}=\frac{\Theta_{1}(v_{1})}{\Theta_{1}(v)+\Theta_{1}(v_{1})},$
$c=\frac{\Theta_{2}(v,v_{2})}{\Theta_{2}(v+v_{1},v_{2})}b,\quad c_{1}=\frac{\Theta_{2}(v_{1},v_{2})}{\Theta_{2}(v+v_{1},v_{2})}b_{1}.$
@@
This establishes that, up to errors bounded by $\nabla^{2}(\kappa,\tilde{N}(\kappa),\tilde{N}(\kappa))$, the $\nabla^{1}$ estimates (and similarly for $\nabla^{2}$) remain valid when replacing one chord in the set $(\kappa)_{II}$ with a chord in the set $(\kappa+1)_{II}$. By iterating this argument, we can successively replace all chords in the set $(\kappa)_{II}$ with chords in the set $(\kappa+1)_{II}$, thereby proving claims (64) and (65).
For the claim (66), consider $(v_{1},v_{2})\in(\kappa+1,\tilde{N}(\kappa+1))_{II}$ and one can find that $v_{3}\in(\kappa+1)_{II}$ satisfies $\Theta_{2}(v_{1},v_{3})\sim q_{\tilde{N}(\kappa)}^{-1}$. Similar to the proof of Lemma 4.4, we have
@@
$d_{1}=-\frac{\Theta_{2}(v_{1},v_{3})}{\Theta_{2}(v_{1},v_{2})},\quad d_{2}=\frac{\Theta_{2}(v_{2},v_{3})}{\Theta_{2}(v_{1},v_{2})},$
@@
$\nabla^{2}(\kappa,\tilde{N},\tilde{N})\leq C_{0}q_{\tilde{N}}^{-\eta-\frac{\varepsilon}{3}}.$
@@
$\tilde{N}_{0}:=\tilde{N}(\kappa_{0}),\quad\varepsilon_{1}:=\frac{\varepsilon}{960A_{\alpha}},\quad C_{5}:=100(A_{\alpha}+1)^{3}C_{0}.$
@@
For each $m>1$, if we assume that for each $\kappa\in[\kappa_{0},\kappa_{0}+m-1]$,
@@
Proof.
By Lemma 5.3 and Lemma 5.4,
@@
$\leq\frac{\varepsilon_{1}}{2}+C_{5}q_{\tilde{N}_{0}}^{-\eta-\frac{\varepsilon}{3}}.$
@@
$\begin{split}\nabla^{1}(\kappa_{0}+m,\tilde{N}(\kappa_{0}+m))\leq\frac{\varepsilon_{1}}{2}+\frac{C_{5}\log_{2}(2A_{\alpha})}{1-\sqrt{2}^{-(\frac{\varepsilon}{3}+\eta)}}\sqrt{2}^{-(\frac{\varepsilon}{3}+\eta)\kappa_{0}}.\end{split}$
@@
(69) $\kappa_{0}>\frac{2\log_{2}\frac{2}{\varepsilon_{1}}\left(\frac{C_{5}\log_{2}(2A_{\alpha})}{1-\sqrt{2}^{-(\frac{\varepsilon}{3}+\eta)}}\right)}{\frac{\varepsilon}{3}+\eta},$
@@
Appendix A Collection of the notions and notations
In the order of appearance in the paper, the main notions and notations are listed below for the reader’s convenience.
  • $A_{\alpha}$ is the upper bound for the partial quotients in the continued fraction expansion of the constant type rotation number $\alpha$.
  • $G_{n}:\mathbb{T}\times\mathbb{T}\to\mathbb{R}$ stands for the sequence of generating functions constructed in the work.
  • $F_{n}:\mathbb{T}\times\mathbb{R}\to\mathbb{T}\times\mathbb{R}$ is the exact area-preserving map generated by $G_{n}$. For a fixed $n\in\mathbb{N}$ we often simply write $F$ in place of $F_{n}$.
  • $\widetilde{F}:\mathbb{R}^{2}\to\mathbb{R}^{2}$ refers to the lift of $F$.
  • $\mathcal{O}$ represents a complete orbit of $F$ with rotation number $\alpha$.
  • $\lambda_{1}(v)$ gives the $\lambda$-length of the chord $v$; $\Theta_{1}(v)$ gives the $\vartheta$-length of the chord $v$. Similarly we define $\lambda_{2}(v_{1},v_{2})$, $\lambda_{4}(v_{1},v_{2},v_{3},v_{4})$ together with $\Theta_{2}$, $\Theta_{4}$, etc.
  • $(\kappa)_{\iota}$ designates the set of Type-$\iota$ $\kappa$-chords, where $\iota\in\{I,II\}$. Accordingly, $(\kappa,r)_{II}$ designates the set of chord pairs $(v_{1},v_{2})$ belonging to this class, and $(\kappa,r,s)_{II}$ with $s\leq r$ the set of four chords of the same class.
  • $n_{\kappa}:=\psi(\kappa)$ denotes, in rough terms, the quantity corresponding to $\kappa$ such that $q_{n_{\kappa}}\sim 2^{\kappa}$ as $\kappa\to+\infty$; the definition of $\psi$ is given in (10) and its inverse is denoted by $\phi$.
  • $\tilde{N}(\kappa)=n_{\kappa}-2\gamma_{0}$ and $\bar{N}(\kappa)=n_{\kappa}+2\gamma_{0}$. These are used to characterize two scales of chord pairs: $(\kappa,\tilde{N}(\kappa))_{II}$ and $(\kappa,\bar{N}(\kappa))_{II}$. A substantial part of the paper is devoted to showing that, under suitable conditions, the estimate on $(\kappa,\tilde{N}(\kappa))_{II}$ can be transferred to $(\kappa,\bar{N}(\kappa))_{II}$. The constant $\gamma_{0}$, which depends only on $\alpha$; its definition can be found in (12).
  • $\Lambda_{\iota}(\kappa)$, where $\iota\in\{I,II\}$, characterise the criteria for the existence of an invariant circle by $F$.
  • $\widetilde{\mathbf{K}}_{N}^{0}(\kappa)$ measures the supremum of the expansion along orbits generated by $F$ whose length does not exceed $q_{N}$; see the precise definition (17). When the length equals exactly $q_{N}$, we write ${\mathbf{K}}_{N}^{0}(\kappa)$.
  • $\widetilde{\mathbf{K}}_{N}^{1}(\kappa,r)$ measures the supremum of the distortion along orbits generated by $F$ whose length is at most $q_{N}$; the exact definition is given in (22). When the length equals exactly $q_{N}$, we write ${\mathbf{K}}_{N}^{1}(\kappa,r)$.
  • $\nabla^{1}(\kappa,r)$ stands for the supremum of the second-order difference quotient of $F$ along $\mathcal{O}$; $\nabla^{2}(\kappa,r,s)$ stands for the supremum of the third-order difference quotient of $F$ along $\mathcal{O}$.
  • $C_{i}$ $(i=0,\ldots,5)$ denotes the positive constant only depending on $\alpha$.